OpenStudy (anonymous):

Factor the expression: x^2 - 4x - 12 a. (x-6)(x-2) b. (x-4)(x-3) c. (x-6)(x+2) d. (x-4)(x+3) What is the solution of the equation? x - 2 = 29 a. -27 b. -31 c. 31 d. 27

OpenStudy (anonymous):

your first question, the answer is D. x * x = x^2, -4 * x = -4x and -4 * 3 = 12 using the distributive property. Don't forget you are multiplying negative and positives. the second question, isolate x by adding 2 to both sides. x=31. by adding 2 on the left, it cancels out to 0 leaving x by itself. What you do to one side you do to the other. So 29+2 equals 31. answer is C.